OpenStudy (anonymous):

Dear All, Please share the formula to find the volume of metal sheet used in a Hollow cylinder ...

OpenStudy (anonymous):

Volume V = π×h×(R² − r²) = π × h × (D² − d²) ⁄ 4 Surface to volume ratio SA:V = 2 × (R − r + h) ⁄ (h × (R − r)) Surface Area (SA), Outer Lateral lS = 2×π×R×h Surface Area (SA), Inner Lateral ls = 2×π×r×h Surface Area (SA), Total Lateral l = lS+ls = (2×π×R×h)+(2×π×r×h)  =  2×π×h×(R+r) Surface Area (SA), Base B = π × (R² − r²) Surface Area (SA), Total A = 2 × B + l  = (2 × π × (R² − r²)) + (2 × π × h × (R + r))  =  2 × π × (R² − r² + h × (R + r))  =  2 × π × (R+r) × (R − r+h) Outer diameter D = 2×R Inner diameter d = 2×r Outer radius R = D ⁄ 2 Inner radius r = d ⁄ 2 Base outer perimeter P = 2×π×R Base inner perimeter p = 2×π×r

OpenStudy (anonymous):

thanks....

OpenStudy (anonymous):

but........I want to information.......about volume of Metal used in a Hollow cylinder.......

OpenStudy (kira_yamato):



OpenStudy (kira_yamato):

What you want is the volume of the metal sheet, essentially the ring. So take the difference between the volume of the external cylinder and the internal one

OpenStudy (anonymous):

Cut the cylinder along its height. It becomes an open sheet. area = pi*d*h vol=pi*d*h*thickness. That is the amount of material is present. Note: 25.4mm=1.0" 10mm=1*10/25.4=0.3936"